Aggregator

Injector work summary

6 days 9 hours ago
Log Number: 4369869Logbook(s): Entry Makers: ywangBody: 

1400-1500 Optimized the injector with beam to ILD:
Figs. 1-4: CW beam from laser A, B, C and D, respectively
Figs. 5-7: B's, C's and D's orbits relative to A's, respectively

ywang

Hall B Halo rates look OK

6 days 9 hours ago
Log Number: 4369868Logbook(s): Entry Makers: mockBody: 

After Injector work, the halo rates in Hall B did not get worse (success!). While optimizing the rest of the injector and removing tail on SLMs.

mock

Testing how injector changes effect Hall B halo rates

6 days 9 hours ago
Log Number: 4369863Logbook(s): Entry Makers: mockBody: 

After Yan made some injector changes. He wanted to check how the halo rates in Hall B would respond. Spoke with the Hall as we told them beam would be off for a longer period of time than they were expecting (sorry!). The Hall agreed to accept some beam to compare halo rates. Starting off at 5 nA and ramping up to 325 nA.

mock

FSD Masking Tool: Successful System Masking for: BeamLossMonitor ***Some node's masks failed to write

6 days 10 hours ago
Log Number: 4369842Logbook(s): Tags: Body: System Masking for: BeamLossMonitor
User: mock
Areas: All

BeamLossMonitor Devices:
ILM0I07
ILM0L01A
ILM0L01B
ILM0L02
ILM2D00A
ILM2D00B
ILM0L04A
ILM0L04B
ILM0L05
ILM0L06A
ILM0L08
ILM0R01
ILM4D00
ILM0R03
ILM0R04
ILM0R06
ILM0R08
ILM0R09
ILM1L01
ILM1L02
ILM6R08
ILM3S00
ILM7S00
ILM9S00
ILM5S10
ILM5E03
ILM7S01
ILM9A01
ILM3E02
ILM5A01
ILM1A04
ILM1A06
ILM1A09
ILM5A10
ILM1A16
ILM1A19
ILM3R10
ILM9R07
ILM1R09
ILM2L03
ILM5R10
ILM5R05
ILM2L01
ILM2T09
ILM8T09<B
mock

Downtime Incident CLOSED: Aluminum Waterflow Alarm

6 days 10 hours ago
Log Number: 4369838Logbook(s): Body: IDDescriptionPeriodCause 14054Title:
Aluminum Waterflow Alarm

Summary:
After repairs were performed, operators began preparing the Hall for beam permit. At which time this alarm came in. Upon investigation, the flow rate appears to read back negative values.
mock

COMMENT ON MPS system maintenance and repairs during run 2025

6 days 10 hours ago
Log Number: 4369822Logbook(s): Body: 

For the upcoming maintenance day we would like to repair some damaged power and signal cables in the North Linac for the BLM's. This would require access to the NL if possible.
We will also install the HV supply to FSD system interface box that we removed earlier for repairs (reference log #4356378.) and has been running on our test stand for 3 weeks with know errors.

sbruhwel

Shift Information

Shift Owl
CEBAF Program Physics!!
MCC Crew Chief Price
MCC Operators Howard, Schoene
Program Deputy Drury
Announcements

Access Status

Recent PSS History

recent pss history graph